OpenStudy (pj754):

"Find the moment about the x-axis of the homogenous region bounded by y = x , y = x + 4." What is it asking for????

zepdrix (zepdrix):

Hmm I dunno :(

OpenStudy (phi):

it sounds like they want they "moment of inertia" of a "plate" that has constant density \(\delta\). \[ I= \int\int_A r^2 \delta \ dA \] where r is the distance from the rotation axis. In this case, is the y value of each point in the region \[ I= \delta \int\int_A y^2\ dA \]

OpenStudy (phi):

But having said all that, the region bounded by y=x and y=x+4 goes on forever, so it's not clear....